Use the chart of comparable companies below to value the stock of Ford: Company Stock Price...

  1. Use the chart of comparable companies below to value the stock of Ford:

Company

Stock Price

Earnings per Share

P/E Ratio

General Motors Co.

40.42

2.35

Toyota Motor Corp.

119.85

11.34

Nissan Motor Co.

17.87

2.15

Ford Motor Co.

1.42

Please Round Answer to 2 Decimal Places.
Answer #1

P/E Ratio stands for Price / Earnings Ratio which is equal to: Stock Price / Earnings Per Share.

The PE ratios for the comparable companies or peers of Ford are as follows:

General Motors Co.: 40.42/2.35 = 17.20

Toyota Motor Co.: 119.85/11.34 = 10.57

Nissan Motor Co.: 17.87/2.15 = 8.31

Ford's value of stock (i.e. stock price can be ascertained by multiplying Ford's P/E Ratio with its earnings. However, the exact P/E ratio of Ford is unknown and has to be estimated. An ideal way of doing it is to take the average P/E ratio of its peers which is :
(17.20 + 10.57 + 8.31) / 3 = 12.02.

If 12.02 is the P/E ratio of Ford, then Ford's stock can be valued at :

P/E Ratio * Earnings Per Share = 12.02*1.42 = 17.07

Note that in the example above, the arithmetic mean of P/E ratio of Ford's peers is used to arrive at the estimated P/E ratio of Ford. However, an investor might look at other averages as well instead of arithmetic mean. One such average could be median.
The median P/E ratio of peers is middlemost value in the array 8.31, 10.57, 17.20 which is 10.57. If median is considered as estimated P/E ratio of Ford, then Ford's stock is valued at 10.57*1.42 = 15.01

In short, Ford's stock deserves a valuation in the range of 15 to 17. But valuations are subjective in the practical world and hence there could be multiple values that could be arrived at.